Let P(x) = x2 - x - 6. Compute P(A) for matrix A = [3, -1and 0 -2]


1
Expert's answer
2021-05-17T08:16:58-0400

Given, "P(x)=x^2-x-6"


"A=\begin{bmatrix} 2&-1\\0&-2\end{bmatrix}"




"P(A)=A^2-A-6I"


   "=\begin{bmatrix} 2&-1\\0&-2\end{bmatrix}\begin{bmatrix} 2&-1\\0&-2\end{bmatrix}-\begin{bmatrix} 2&-1\\0&-2\end{bmatrix}-6\begin{bmatrix}1&0\\0&1\end{bmatrix}
\\[9pt]


= \begin{bmatrix} 9&-1\\0&4\end{bmatrix} +\begin{bmatrix} -3&1\\0&2\end{bmatrix} -\begin{bmatrix} 6&0\\0&6\end{bmatrix} 
\\[9pt]


=\begin{bmatrix} 0&0\\0&0\end{bmatrix}"